OpenStudy (anonymous):

Which set of line segments with the given lengths form a right triangle? 5, 12, 17 7, 11, 15 8, 15, 17 14, 16, 20

OpenStudy (anonymous):

8, 15, 17

OpenStudy (anonymous):

8,15,17 8^2+15^2=17^2

OpenStudy (anonymous):

8^2+15^2=17^2